### v3

@@ 2.5 Thermodynamic and kinetic properties
(b) Diffusivities
Figure 5: (a) Comparison of experimental activity coefficient derivatives with predictions using Kirkwood-Buff theory. (b) Comparison of the computed diffusivities with experimental values[67] for water and both ions at two concentrations and infinite dilution.
The reliability of the ion-ion RDFs can be validated by comparing them with thermodynamic properties. Specifically, the activity coefficients as a function of concentration have been shown to be highly sensitive to the strength of ion-ion interactions. The accurate prediction of these quantities provided the original validation of Debye-Hückel theory. The ability to determine these properties accurately is also of immense practical importance for chemical engineering, where they help determine a range of properties such as equilibrium constants, solubilities, and reaction rates.
Here, we use Kirkwood-Buff theory to compute a derivative of the log activities, showing good experimental agreement as shown in Figure 5(a). The stronger pairing of KCl compared to LiCl and the slightly weaker pairing of LiBr are reproduced. The stronger pairing of potassium to chloride is a classical example of a counterintuitive ‘specific ion effect’ induced by the solvent, which reverses the expected pairing strength in vacuum. In contrast, the stronger pairing of lithium chloride with bromide does follow the expected behaviour in vacuum and is reproduced. Reproducing these binding strengths is key to understanding many important phenomena such as the effect of ions on protein stability, i.e., the so-called ‘Hofmeister effect’.[68]
We have also computed the activities at 1.3 M for LiCl and 6.6 M for KCl to demonstrate the generalizability to different concentrations. This requires correctly reproducing the change in Debye screening length as a function of concentration. The fact that NNP-MD can reproduce this effect, despite using training data from a single concentration suggests that the NNP is learning a correct representation of the water-water interactions rather than merely interpolating on the training dataset.
Note that the KCl experimental values rely on extrapolations of the experimental data, as this is above the experimental solubility point. This highlights the capability of this method to obtain or confirm experimentally inaccessible data.
Diffusivities of Li+ and Cl- ions, as well as water molecules, are also in good agreement with experiment, as shown in Figure 5(b). This is particularly impressive as kinetic properties depend on accurately assessing barrier heights in the potential energy surface. These will be poorly represented in the training dataset as it is extracted from equilibrium molecular dynamics.
2.6 Continuum solvent coarse-graining
(a) LiCl at 2.5 M.
@@ 2.6 Continuum solvent coarse-graining
Figure 6: Comparison of all-atom RDFs with coarse-grained RDFs for (a) LiCl at 2.5 M and (b) KCl at 6.6 M.
The all-atom NNP-MD is much faster than FPMD and capable of simulating experimentally relevant timescales, but the computational cost is still non-trivial and more expensive than most classical molecular dynamics (CMD) approaches. To further lower the computational cost of our method, we build a coarse-grained model of the electrolyte solutions. Specifically, we integrate out the solvent degrees of freedom, resulting in a continuum or implicit solvent model. To do this, we use the NNP to learn the potential of mean force, which is a free energy surface. The problem of learning a free energy surface is very similar to learning the full all-atom potential energy surface and thus also benefits from equivariance, as has recently been demonstrated.[23] As before, we compute the long-range electrostatic interactions separately using Coulomb’s law.
In practice, to do this for LiCl, we extract the coordinates and forces for the ions alone for 24,000 frames extracted from a 2.4 ns NNP-MD all-atom simulation. A larger dataset is required to sufficiently converge the averaging over the solvent degrees of freedom. It would not be feasible to generate such a large training dataset with FPMD directly, but it is straightforward with NNP-MD. The coarse-grained NNP requires many fewer weights and trains very quickly in comparison to the all-atom NNP due to the much simpler energy surface.
The coarse-grained MD can accurately reproduce the RDFs from the all-atom MD, as shown in Figure 6(a).
Additionally, to further reduce the training data requirements for the coarse-grained model we ran the KCl all-atom NNP (trained at 2.5 M) at 6.6 M and used 7500 frames from that to train a coarse-grained model which is able to reproduce the all-atom RDFs highly accurately, as shown in Figure 6(b). Surprisingly, the all-atom NNP trained at 2.5 M KCl is also capable of reproducing the 6.6 M RDFs very accurately.
@@
This demonstrates the capability to generalise to higher concentration. This is surprising given that higher-order many-body effects are likely to occur at higher concentration that have not been observed at the lower concentrations. This constitutes further evidence of the correct physical interactions being learnt.
The coarse-grained NNP-MD is orders of magnitude faster than the all-atom NNP and requires trivial computational resources (tens of CPU hours) to fully converge. This hierarchical layering of NNPs where a coarse-grained NNP is trained on data from all-atom NNP-MD is promising as a general approach to the long-standing challenge of connecting scales in molecular simulation given its simplicity. [69, 70]
@@
To further validate the reliability of this approach, we also trained a coarse-grained NNP on a classical force field for NaCl and demonstrated that RDFs can be reproduced at different concentrations reliably, as shown in Figure 10. We did not train on first principles data for NaCl, as the standard pseudopotential introduces some noise in the forces on the Na ion. This can be corrected with the Gaussian and Augmented Plane Wave (GAPW) method, but this has not been implemented with DC-DFT yet.[45] Additionally, classical force fields for NaCl have already been carefully parameterised to experiment and can do a reasonable job computing properties such as activity coefficients.[71]
(a) LiCl PMFs computed from the CG model.
(b) Classical NaCl short range PMFs.
Figure 7: (a) Infinite dilution PMFs for LiCl computed with the coarse-grained model trained at 2.5 M. (b) Infinite dilution PMFs for the NaCl classical force-field computed with the coarse-grained model at various concentrations compared with direct metadynamics and thermodynamic integration of a single pair.
@@
One key property of solutions that has, to date, been difficult to compute is the infinite dilution potential of mean force (PMF), also called the pair potential. This corresponds to the free energy change as two solutes move apart in water at infinite dilution, i.e., infinite dilution pairing free energy. Previously, computation of this quantity required enhanced sampling techniques, such as umbrella sampling with large box sizes. However, our method allows us to trivially extract this quantity from the coarse-grained NNP by simply computing the interaction energy between the ion pairs in isolation. These infinite dilution PMFs are shown in Figure 7(a) for LiCl, with the long-range Coulomb term removed. We validate the reliability of this procedure by showing that the PMF calculated in this way reproduces the true infinite dilution PMF calculated for a classical force field using metadynamics and thermodynamic integration, as shown in Figure 7(b).
@@
Figure 8: (a) NaCl nucleation and crystal growth observed with a coarse-grained model trained on solution-phase classical MD simulation. (b) KCl dissolution with a coarse-grained model trained only on first principles solution-phase data. The snapshots are evenly spaced.
Most remarkably, while running the coarse-grained NNP trained on the classical NaCl force field, we observed stable crystal nucleation and growth (Fig. 8(a)), despite the fact that this model was only trained on solution-phase simulations. While we never observed this behaviour in the all-atom MD despite very long simulations, this is actually the physically correct behaviour as this classical force field does have a relatively low solubility.[72] The lattice spacing is too large at 6.21 Å  compared to the experimental value of 5.63 Å  but the correct FCC-structured crystal is formed. This indicates that it should be possible to study complex phenomena, such as crystal nucleation and growth, with a coarse-grained model.
We did not observe spontaneous nucleation with the coarse-grained KCl/LiCl force fields, likely due to their much higher solubility. However, we did initialise a simulation with a KCl crystal and observed continuous slow dissolution into the liquid starting from the corners while generally maintaining the correct crystalline structure, including the correct lattice parameter of 6.3 Å. This was followed by a final rapid dissolution step (Fig. 8(b)). Such behaviour has also recently been observed for NaCl in all-atom NNP-MD simulation.[73] The significant difference, in this case, is that our simulations with the coarse-grained model are run on a single CPU within a day with no requirements for crystal-phase training data. A movie of the KCl crystal dissolution and NaCl nucleation is available online.
@@ 3 Conclusion and future work
In summary, we have demonstrated the accurate prediction of the structural, kinetic, and thermodynamic properties of aqueous LiCl, KCl, and LiBr electrolyte solutions. The prediction of these properties, particularly activity coefficients, has been a key goal of physical chemistry for over a century. In the process of doing so, we also discovered the formation of lithium dimers and nearly identical anion-anion interactions of chloride and bromide. Our approach combines equivariant NNPs and DC-DFT to achieve this predictive ability with reasonably of training data and relatively few computational resources.
In addition, we have demonstrated the ability to recursively train an NNP on its own output to build coarse-grained continuum solvent NNPs capable of reproducing all-atom RDFs with further reduced computational demand. These coarse-grained models are capable of reproducing crystal phase behaviour as well as infinite dilution pairing free energies despite being trained only on moderate concentration solution.
This approach should be applicable to a wide range of electrolyte solutions. Although, generalising to more complex systems may require more advanced techniques for generating the datasets, such as meta-dynamics[74] and active learning[40, 44] to further improve stability. This should make it possible to build a database of properties of electrolytes across a much wider range of conditions and compositions than currently exists, a task of critical industrial importance. A key focus should be on important electrolytes that are particularly difficult to characterise experimentally, such as pure lithium bicarbonate, which immediately speciates into a mixture of carbonates in solution, meaning even its most basic properties have not been directly measured. Additionally, properties under high-temperature and high-pressure conditions, where experimental data is hard to obtain, should be a key target.
@@ 5.1 FPMD (CP2K)
We used Born-Oppenheimer ab initio molecular dynamics simulations within a constant volume NVT (300 K) ensemble with periodic boundary conditions. The CP2K simulation suite, containing the QuickStep module for the DFT calculations [59, 60], was used with a 0.5 fs time step. We used a double $\zeta$ basis set that has been optimized for the condensed phase[75] in conjunction with GTH pseudopotentials [76] optimised for SCAN[77, 78] and a 1200 Ry cutoff.[79, 80] A slightly smaller basis set (DZVP-MOLOPT-SCAN-GTH) was used for LiBr to test the sensitivity of the lithium ion pairing to the basis set size. A Nosé-Hoover thermostat was attached to every degree of freedom to ensure equilibration. [81]
An $\approx$ 10 ps simulation was run, consisting of 4 cations ions, 4 anions ions, and 80 water molecules for each electrolyte. Cells with fixed dimensions of 13.873 Å3, 14.02233 Å3, and 13.7 3 Å3 cell were used for LiCl, KCl and LiBr respectively. Corresponding to an electrolyte concentration of 2.5 M. The cell size was adjusted to match the experimental density at this concentration. The initial simulation used the r2SCAN DFA.[52]
@@ 5.2 NNP fitting (NequIP)
The training dataset was generated as follows: For LiCl, KCl, and LiBr, 458, 492, and 322 frames were sampled directly from an $\approx$ 10 ps FPMD run using r2SCAN. These datasets were used to generate $\approx$ 1 nanosecond of NNP-MD data at the FPMD box size. 197, 207, and 201 (30%) frames were then sampled from the 1 ns NNP-MD run to improve the diversity of training data for the more sophisticated NNP, resulting in a total of 655,699, and 523 frames.
@@
As our NNP only has access to local information (short range, $<5$[19] )the long-range electrostatic ion-ion interactions were removed from the forces and energies prior to training. These were computed using a dielectrically screened Coulomb interaction.[83] They were then added back in during all NNP-MD simulations. These were calculated with LAMMPS by placing appropriately screened charges on the ions to reproduce dielectric screening and were computed with the particle-particle particle-mesh method.[84] For KCl and LiBr, a lower dielectric of 50 was used to better describe the high concentration regime to account for the decrease with ion concentration. We tested this for LiCl to confirm did not have a significant effect at 2.5 M.
@@
The same hyperparameters were used for the all-atom NNPs. More specifically, 100:1 weighting on forces vs. energies was used in the default loss function.[14] We decreased the initial learning rate of 0.01 by a decay factor of 0.5 whenever the validation RMSE in the forces did not see an improvement for five epochs. Training was stopped when the learning rate became smaller than 10-5. The model with the lowest validation error was used for simulations. A radial cutoff distance of 5 Å was used. Three layers of interaction blocks were used with the maximum $l$ set to 2, each with 16 features. Only even parity was used. Invariant neurons for the radial network was set to 32. All the other parameters were set to the defaults. An 80:20 training-validation split was used throughout.
@@
To train the NNP for the coarse-grained MD, we found that reducing the number of parameters was important to provide stability. Two layers of interaction blocks were used with the maximum $l$ set to 1 each with 8 features. Only even parity was used. Invariant neurons for the radial network were set to 8. All the other parameters were set to the defaults. The radial cutoff was also extended to 10 Å to provide longer-range interactions.
@@
Figure 9 shows the learning curves. The RMSE on the validation set for the all-atom NNPs was between 9-12 meV/Å for the forces and 0.1-0.14 meV for the energies.
@@
(a) All-atom LiCl NNP
(b) coarse-grained LiCl NNP
(c) All-atom KCl NNP
(d) coarse-grained KCl NNP
(e) coarse-grained LiBr NNP
Figure 9: Learning curves for all-atom and coarse-grained NNPs.
To test the KCl at high concentrations, 261 frames were extracted from the 6.6 M run and resampled with DC-r2SCAN. These were used to train a new all-atom neural network potential, which showed good agreement with the results of the lower concentration model.

### v4

@@ 2.3 Ion pairing: prediction and validation
(c) LiCl PMFs computed from the CG model.
(d) Classical NaCl short range PMFs.
Figure 4: (a) Comparison of experimental activity coefficient derivatives with predictions using Kirkwood-Buff theory and RDFs from the all atom NNP-MD simulations. (b) Comparison of all-atom RDFs with coarse-grained RDFs for KCl at 6.6 M. (c) Infinite dilution PMFs for LiCl computed with the coarse-grained model trained at 2.5 M using DFT data. (d) Short ranged infinite dilution potentials of mean force (srPMFs), i.e., free energies, for the NaCl classical force-field computed with the coarse-grained model at various concentrations compared with direct calculation using enhanced sampling. The long range Coulomb component has been removed and thermal units on the y axis are used for (c) and (d). Metadyn. refers to metadynamics, a direct method of computing PMFs.
We have also predicted the activity coefficient derivatives at 1.3 M for LiCl and 6.6 M for KCl to demonstrate the generalisability to different concentrations not included in the training data. Activity coefficients describe the non-ideal free energy of ions in solution and are critically important for a wide range of applications. Reproducing these requires correctly reproducing the change in Debye screening length as a function of concentration. (See Figure S6). Debye screening refers to the fact that electrostatic interactions become increasingly short range in electrolytes as the concentration increases due to the cloud of counter charge around an ion. The fact that NNP-MD can reproduce this effect, despite using training data from a single concentration (2.5 M) indicates that the NNP is learning a correct representation of the water-water interactions. That it is accurate at higher concentrations (6.6 M) is particularly surprising given that higher-order many-body effects are likely to occur here that have not been observed at the lower concentrations. Note that the KCl experimental values rely on extrapolations of the experimental data, as this is above the experimental solubility point. (Figure 4(a)) This highlights the capability of this method to obtain experimentally inaccessible data.
@@
This demonstrates the capability to generalise to higher concentration. This constitutes further evidence of the correct physical interactions are being learnt.
Diffusivities of Li+ and Cl- ions, as well as water molecules, are also in reasonable agreement with experiment, as shown in Figure S4. This is particularly impressive as kinetic properties, such as diffusivities, depend on accurate barrier heights in the potential energy surface. Structures at the barrier heights will be underrepresented in the training dataset as they are extracted from equilibrium molecular dynamics.
2.4 Coarse-graining: prediction and validation
The all-atom NNP-MD is much faster than FPMD and capable of simulating experimentally relevant timescales, but the computational cost is still non-trivial and more expensive than most classical molecular dynamics (CMD) approaches. To further lower the computational cost of our method, we build a coarse-grained/continuum-solvent NNP of the electrolyte solutions. Specifically, we integrate out the solvent degrees of freedom, resulting in a continuum or implicit solvent model. To do this, we train an coarge grained NNP to learn the forces on the ions averaged over the equilibrium water configurations. The training data comes all atom MD simulations (either classical or NNP). This corresponds to learning a free energy surface, also referred to as a potential of mean force. Learning free energies is a very similar task to learning the full all-atom potential energy surface and thus also benefits from equivariance, as has recently been demonstrated for pure water.[23] As before, we compute the long-range electrostatic interactions separately using Coulomb’s law. The coarse-grained MD can accurately reproduce the RDFs from the all-atom MD, as shown by the close match between dashed and solid lines in Figure 4(b). Good agreement is also observed for LiCl (see Figure S6).
The coarse-grained NNP-MD is much faster than the all-atom NNP, producing fully converged RDFs within a day on a single CPU, this is due to the faster inference time (300 ps per day on a single CPU vs 200 ps per day on a GPU) and due to the fact that coarse-graining accelerates the dynamics. This hierarchical layering of NNPs where a coarse-grained NNP is trained on data from all-atom NNP-MD is promising as a general approach to the long-standing challenge of connecting scales in molecular simulation given its simplicity. [70, 71] A comparison of the accuracy of the coarse grained and all atom models is difficult as the ground truth PMFs that the coarse grained model are predicting are unknown. It has much higher loss than the all-atom model, but this is due to the inherent noise from the neglect of solvent degrees of freedom, rather than inherent inaccuracy in the model.
2.5 Infinite dilution
@@
To further validate the reliability of this approach, we also trained a coarse-grained/continuum-solvent NNP on a classical all atom force field for NaCl and demonstrated that RDFs can be reproduced at different concentrations reliably, as shown in Figure S8. We did not train on first principles data for NaCl, as the standard pseudopotential introduces some noise in the forces on the Na ion. This can be corrected with the Gaussian and Augmented Plane Wave (GAPW) method,[45] but this has not been implemented with DC-DFT yet. Additionally, classical force fields for NaCl have already been carefully parameterised to experiment and can do a reasonable job computing properties such as activity coefficients.[72]
@@
One key property of solutions that has, to date, been difficult to compute is the infinite dilution potential of mean force (PMF), also called the pair potential. This corresponds to the free energy change as two solutes move apart in water at infinite dilution, i.e., the infinite dilution pairing free energy. Previously, computation of this quantity required enhanced sampling techniques, such as umbrella sampling with large box sizes. However, our method allows us to trivially extract this quantity from the coarse-grained NNP by simply computing the interaction energy between the ion pairs in isolation. These infinite dilution PMFs are shown in Figure 4(c) for LiCl, with the long-range Coulomb term removed. We validate the reliability of this procedure by showing that the PMF calculated in this way reproduces the ground truth infinite dilution PMF for this classical force field as shown in Figure 4(d). This was computed using metadynamics and thermodynamic integration. See Ref  for details. This indicates that coarse-grained/continuum-solvent NNPs are a promising tool for the rapid calculation of intrinsic binding free energies of species in solution a problem of ubiquitous importance throughout molecular science.
2.6 Crystallisation
@@
Figure 5: (a) NaCl nucleation and crystal growth observed with a coarse-grained NNP-MD simulation trained on solution-phase classical all atom MD simulation. (b) KCl dissolution with a coarse-grained NNP-MD simulation trained on an all atom NNP-MD simulation, which was trained on DFT data. The snapshots are evenly spaced.
Most remarkably, while running the coarse-grained NNP trained on the classical NaCl force field, we observed stable crystal nucleation and growth (Fig. 5(a)), despite the fact that this model was only trained on solution-phase simulations. While we never observed this behaviour in the all-atom MD despite long simulations, this is actually the physically correct behaviour as the solubility predicted by this classical force field is relatively low.[73] The lattice spacing, i.e., the edge length of the unit cell, is too large at 6.21 Å  compared to the experimental value of 5.63 Å  but the correct FCC-structured crystal is formed. This indicates that it should be possible to study complex phenomena, such as crystal nucleation and growth, with a coarse-grained/continuum-solvent NNP.
We did not observe spontaneous nucleation with the coarse-grained KCl/LiCl NNP, likely due to their much higher solubility. However, we did initialise a simulation with a KCl crystal and observed continuous slow dissolution into the liquid starting from the corners while generally maintaining the correct crystalline structure, including the correct lattice parameter of 6.3 Å. This was followed by a final rapid dissolution step (Fig. 5(b)). Such behaviour has also recently been observed for NaCl in all-atom NNP-MD simulation.[74] The significant difference, in this case, is that our coarse grained NNP simulations are run on a single CPU within a day with no requirements for crystal-phase training data.
We also simulated a slab of KCl crystal in periodic boundary conditions with a vacuum region for 1 ns. The vacuum region quickly become saturated with liquid phase KCl ions after which the stable crystal structure of KCl was maintained, indicating that the crystal was not inherently unstable. A movie of the KCl crystal dissolution and NaCl nucleation is available online.
The use of the output of the all atom neural network potential simulations to train the coarse grain simulations may cause concerns around model collapse, a known issue with training on synthetic data. However, this is not a concern in our case as we are not training on the raw output of a model but instead on molecular simulation data generated with a conservative force field. The coarse graining is therefore a mathematically rigorous process and unlikely to be subject to the same issues. The extent to which this process can be repeated to access even larger scales however, remains an open interesting question.
@@ 3 Conclusion and future work
In summary, we have demonstrated the accurate prediction of the structural, kinetic, and thermodynamic properties of aqueous LiCl, KCl, and LiBr electrolyte solutions. The prediction of these properties, particularly activity coefficients, has been a key goal of physical chemistry for over a century. In the process of doing so, we also discovered the formation of lithium dimers and nearly identical anion-anion interactions of chloride and bromide.
In addition, we have demonstrated the ability to recursively train an NNP on its own output to build coarse-grained/continuum-solvent NNPs capable of reproducing all-atom RDFs with further reduced computational demand. These coarse-grained models are capable of reproducing crystal phase behaviour as well as infinite dilution pairing free energies despite being trained only on moderate concentration solution.
Importantly, our approach combines equivariant NNPs and DC-DFT to achieve this predictive ability with reasonable training data requirements and moderate computational resources. This means it is feasible to scale this approach to many different electrolyte solutions, enabling the construction of a database of properties of electrolytes across a much wider range of conditions and compositions than currently exists, a task of critical industrial importance. A key focus should be on important electrolytes and properties that are particularly difficult to characterise experimentally, as well as high-temperature and high-pressure conditions, where experimental data is difficult to obtain.
@@ 5.1 FPMD (CP2K)
We used Born-Oppenheimer first principles molecular dynamics (FPMD) simulations within a constant volume NVT (300 K) ensemble with periodic boundary conditions. The CP2K simulation suite, containing the QuickStep module for the DC-DFT calculations [60, 61], was used with a 0.5 fs time step. We used a double $\zeta$ basis set that has been optimized for the condensed phase[75] in conjunction with GTH pseudopotentials [76] optimised for SCAN[77, 78] and a 1200 Ry cutoff.[79, 80] A slightly smaller basis set (DZVP-MOLOPT-SCAN-GTH) was used for LiBr to test the sensitivity of the lithium ion pairing to the basis set size. A Nosé-Hoover thermostat was attached to every degree of freedom to ensure equilibration. [81]
An $\approx$ 10 ps FPMD simulation was run, consisting of 4 cations ions, 4 anions ions, and 80 water molecules for each electrolyte. Cells with fixed dimensions of 13.93 Å3, 14.03 Å3, and 13.7 3 Å3 cell were used for LiCl, KCl and LiBr respectively, corresponding to an electrolyte concentration of 2.5 M. The cell size was adjusted to match the experimental density at this concentration. The initial FPMD simulation used the r2SCAN DFA.[52]
@@ 5.2 NNP fitting (NequIP)
The NequIP package was used for the NNPs.[18] The training dataset was generated as follows: For LiCl, KCl, and LiBr, 458, 492, and 322 frames were evenly sampled directly from an $\approx$ 10 ps FPMD run using r2SCAN. These datasets were used to generate $\approx$ 1 nanosecond of all atom NNP-MD data at the FPMD box size. 197, 207, and 201 frames were then evenly sampled from the 1 ns NNP-MD run and force and energies were computed with DC-DFT and added to the training data to improve the diversity for the a second NNP, resulting in a total of 655, 699, and 523 frames for LiCl, KCl and LiBr respectively.
@@
As our NNP only has access to local information (short range, $<5$[14] )the long-range electrostatic ion-ion interactions were removed from the forces and energies prior to training. These were computed using a dielectrically screened Coulomb interaction.[83] They were then added back in during both all atom and coarse grained NNP-MD simulations. These were calculated with LAMMPS by placing appropriately screened charges on the ions to reproduce dielectric screening and were computed with the particle-particle particle-mesh method.[84] Charges of $+1$ for the cations and $-1$ for the anions were used. For KCl and LiBr, a lower dielectric of 50 was used to better describe the high concentration regime to account for the decrease with ion concentration. Whereas for LiCl 78.3 was used, however, we tested LiCl with a dielectric constant of 50 and confirmed that there was no significant difference at 2.5 M.
@@
The same hyper-parameters were used for the all-atom NNPs. More specifically, 100:1 weighting on forces vs. energies was used in the default loss function.[18] We decreased the initial learning rate of 0.01 by a decay factor of 0.5 whenever the validation RMSE in the forces did not see an improvement for five epochs. Training was stopped when the learning rate became smaller than 10-5. The model with the lowest validation error was used for simulations. A radial cutoff distance of 5 Å was used. Three layers of interaction blocks were used with the maximum $l$ set to 2, each with 16 features. Only even parity was used. Invariant neurons for the radial network was set to 32. All the other parameters were set to the defaults. An 80:20 training-validation split was used throughout. The validation loss generally appears higher than the training loss perhaps indicating under-fitting, additional optimization of the hyper-parameters is likely possible.
@@
To train the NNP for the coarse-grained MD with LiCl, we extract the coordinates and forces for the ions alone for 24,000 frames extracted from a 2.4 ns NNP-MD all-atom simulation. A larger dataset is required to sufficiently converge the averaging over the solvent degrees of freedom. It would not be feasible to generate such a large training dataset with FPMD directly, but it is straightforward with NNP-MD. The coarse-grained NNP requires many fewer weights and trains very quickly in comparison to the all-atom NNP due to the much simpler energy surface, we found that reducing the number of parameters was important to provide stability, meaning less physical inaccuracies in the simulation. Two layers of interaction blocks were used with the maximum $l$ set to 1 each with 8 features. Only even parity was used. Invariant neurons for the radial network were set to 8. All the other parameters were set to the defaults. The radial cutoff was also extended to 10 Å to provide longer-range interactions.
@@
Figure S3 shows the learning curves. The RMSE on the validation set for the all-atom NNPs was between 9-12 meV/Å for the forces and 0.1-0.14 meV for the energies.
Figure S5 shows a histogram of errors for the all atom LiCl NNP, the linear correlation plot is shown in Figure 2(a).
@@
To test the all atom NNP could simulated KCl at high concentrations, 261 frames were extracted from the 6.6 M run and resampled with DC-r2SCAN. These were used to train an additional all-atom neural network potential, which showed good agreement with the results with the model trained on data at 2.5 M. (See Figure S7)